a square block weighing 1.1kn and 250 mm on an edge slides down an incline on a film of foil 6.0μm thick. assuming a linear velocity profile in the oil and neglecting air resistance, what is the terminal velocity of the block? the viscosity of oil is 7 Mpa-s angle of inclination to 20°.

1
Expert's answer
2022-02-17T12:17:02-0500

Terminal velocity = "\frac{(250 mm)2}{1.1kn}""\times" "\frac{6.0μm}{ 7 Mpa-s}"


Terminal velocity = 30.9592m/s